given that c=200+0.8(y-t).
a. what is the equilibrium level of national income?
b. derive the savings function and find the level of savings.
1
Expert's answer
2014-10-10T11:51:02-0400
c = 200 + 0.8(y - t)
a. the equilibrium level of national income will be:
c = 200 + 0.8y - 0.8t, so
0.8y = c + 0.8t - 200
y = 1.25c + t - 250
b. derive the savings function and find the level of savings.
s = y - c - t = y - 200 - 0.8y + 0.8t - t = 0.2y - 0.2t - 200 = 0.2(y - t) - 200